OpenStudy (abmon98):

(1+cos(theta))(1/sin(theta)-1/tan(theta)=sin(theta)

OpenStudy (***[isuru]***):

u mean \[[ (1 + \cos \theta ) \times \frac{ 1 }{ \sin \theta } ] - \frac{ 1 }{ \tan \theta } = \sin \theta \] ??

OpenStudy (***[isuru]***):

or \[[ 1 + \cos \theta ] [ ( \frac{ 1 }{ \sin \theta } - \frac{ 1 }{ \tan \theta } ) ] = \sin \theta\] i think it should be this one

OpenStudy (abmon98):

yes this one

OpenStudy (***[isuru]***):

first step write \[\frac{ 1 }{ \tan \theta }\] using sin and cos ... what will u get ?

OpenStudy (***[isuru]***):

\[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\] sooo.... \[\frac{ 1 }{ \tan \theta } = ? \]

OpenStudy (anonymous):

[1+cosθ][(1/sinθ−1/tanθ)]=sinθ [1 + cosθ] [1/sinθ -cosθ/sinθ] [1 + cosθ][1-cosθ/sinθ] 1 -cos^2 θ/sinθ sin^2 θ/sinθ sinθ